OpenStudy (anonymous):

For this question -2x + y = 6, find the slope and y intercept, and write the equation in slope intercept form.

OpenStudy (unanimoose):

The y intercept is the By. So what is the value that is placed in front of y?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Is this right

OpenStudy (unanimoose):

Woops, I think i did something wrong. Hold on.

OpenStudy (anonymous):

ok

OpenStudy (unanimoose):

Ok yeah we have to convert this equation to slope-intercept form. Do you know what slope intercept form is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=mx +b

OpenStudy (unanimoose):

Yup, do you know how we will be converting -2x + y = 6 into slope-intercept form?

OpenStudy (anonymous):

Do have to subtract from the equation

OpenStudy (anonymous):

I mean the y

OpenStudy (unanimoose):

Well we have to isolate it, as in take -2x and put it n the other side to leave the y by itself and match it with y = mx + b. Do you know what the first step is?

OpenStudy (anonymous):

Add -2x to each side of the equation

OpenStudy (unanimoose):

Yes.

OpenStudy (anonymous):

is this the first step

OpenStudy (unanimoose):

Yup.

OpenStudy (anonymous):

y= 2x + 6

OpenStudy (unanimoose):

Yup! Now we know that is the slope and b is the y intercept, so with this can you answer your initial question?

OpenStudy (unanimoose):

Pardon, m is the slope.

OpenStudy (anonymous):

ok

OpenStudy (unanimoose):

So what is the slope and the y-intercept?

OpenStudy (anonymous):

That was the part i was stuck on

OpenStudy (unanimoose):

Well let's look. y = mx + b y = 2x + 6 m = ? b = ?

OpenStudy (anonymous):

M=2 b=6

OpenStudy (anonymous):

is this right

OpenStudy (unanimoose):

Yes!

OpenStudy (unanimoose):

So M = slope and b = y-intercept, which should give you your answer.

OpenStudy (anonymous):

y= 2x +6

OpenStudy (unanimoose):

Yes.

OpenStudy (anonymous):

Thanks for your help